# Molecule Language Model with Augmented Pairs and Expertise Transfer

### v1

@@ Abstract
Understanding the molecules and their textual descriptions via molecule language models (MoLM) recently got a surge of interest among researchers. However, unique challenges exist in the field of MoLM due to 1) a limited amount of molecule-text paired data and 2) missing expertise that occurred due to the specialized areas of focus among the experts. To this end, we propose AMOLE, which 1) augments molecule-text pairs with structural similarity preserving loss, and 2) transfers the expertise between the molecules. Extensive experiments on various downstream tasks demonstrate the superiority of AMOLE in comprehending molecules and their descriptions, highlighting its potential for application in real-world drug discovery. The source code for AMOLE is available at  https://github.com/Namkyeong/AMOLE.
1 Introduction
@@
Recently, Language Models (LMs) have gained traction in molecular science, fueled by the abundance of literature in this domain. Notably, the conventional method of representing molecules as strings, such as SMILES strings (Weininger, 1988), facilitates the integration of two different modalities, i.e., text and molecules, into a single LM. Then, following the masked language modeling (Devlin et al., 2018), Zeng et al. (2022) train the model on masked SMILES and text from the scientific literature. Moreover, inspired by T5 models (Raffel et al., 2020), training the Molecule Language Model (MoLM) with multiple tasks and fine-tuning the downstream task has been proposed (Edwards et al., 2022; Pei et al., 2023; Christofidellis et al., 2023). However, all these models require molecules to be represented in a string format, like SMILES, to be understood by language models.
@@
As an alternative approach, one can treat molecules and language as multiple modalities, following the recent success of the vision language model (VLM). For instance, several recent works (Edwards et al., 2021; Liu et al., 2023a; Seidl et al., 2023) have proposed using separate encoders for each modality and training the paired modalities to be similar in the representation space using contrastive learning, inspired by the CLIP model (Radford et al., 2021). These models are effective in capturing complementary information between different modalities by learning a joint representation space that can be utilized for various downstream tasks such as cross-modal retrieval and molecular property prediction.
@@
Figure 1: (a) Rivastigmine’s textual descriptions from various experts. (b) The majority of molecules in the PubChem database have only one description provided by an expert.
@@
Despite promising early strides, the progress of MoLM lags far behind its VLM counterparts due to the scarcity of molecule-text paired data, both in quantity and expertise. First, in terms of quantity, the VLM community largely follows the viewpoint that scale is everything, as image-text pairs are widely available on the web (Wang et al., 2023). Pre-training VLM on the massive amounts of crawled image-text pairs, ranging from tens of thousands to billions, consistently leads to significant performance gains in various downstream tasks (Jia et al., 2021; Yu et al., 2022). In contrast, MoLM faces a bottleneck due to the limited amount of molecule-text paired data available, which demands costly domain knowledge and significant time investment for wet lab experiments.
@@
Moreover, each molecule has various descriptions provided by different experts, each focusing on its unique areas of expertise. For example, in the case of a molecule named Rivastigmine (Figure 1 (a)), FDA explains its function as a cholinesterase inhibitor, whereas LiverTox (Hoofnagle, 2013) focuses on its effectiveness on Alzheimer’s disease. However, due to the specialized areas of the experts, they typically restrict their knowledge to a selective group of molecules, resulting in numerous molecules having missing expertise across various experts. Specifically, our examination of the largest molecule database, PubChem Kim et al. (2021), as illustrated in Figure 1 (b), reveals that only 27K out of 299K molecules have descriptions from multiple experts. Consequently, the remaining 272K molecules are each documented with a single expert description, which may lack comprehensive expertise about the molecule.
@@
In this paper, we propose AMOLE which addresses the unique challenges faced in MoLM by Augmenting MOLecule-text pair and transferring Expertise between the molecules. To overcome the lack of molecule-text paired data, we utilize the idea that molecules with similar structures have similar properties, which is grounded in the well-established biochemical principle (Martin et al., 2002). Specifically, we propose to augment molecule-text pairs by sharing descriptions among structurally similar molecules. However, this approach can lead to false positives, as the descriptions may not be specifically written for structurally similar molecules. To address this issue, we introduce a novel loss function that preserves structural similarity, guiding the model to align the augmented molecule-text pairs more closely based on the structural similarity of the molecules.
@@
We make the following contributions:
@@
2 Related Works
@@
MoLM with a Single Language Model. Thanks to the wealth of literature and the traditional string-based representation of molecules, such as SMILES, LMs have been applied to the domain of molecular science. Drawing inspiration from the masked language model approach used in BERT training (Devlin et al., 2018), KV-PLM (Zeng et al., 2022) proposes to train LMs by reconstructing masked SMILES and textual data. Moreover, MolT5 (Edwards et al., 2022) proposes to pre-train the LMs with the “replace corrupted spans” objective (Raffel et al., 2020) on both SMILES string and textual data, followed by fine-tuning for tasks such as molecule captioning and generation. Pei et al. (2023) and Christofidellis et al. (2023) extend MolT5 with various pre-training tasks, such as protein FASTA reconstruction and chemical reaction prediction. However, these models rely on string-based representations of molecules, e.g., SMILES, which are broadly recognized for their lack of topology awareness (Rong et al., 2020). Furthermore, merging two modalities into a single model prevents the adoption of existing pre-trained models tailored for each modality (Liu et al., 2023a).
@@
MoLM with Multi-Modal Contrastive Learning. Inspired by the recent success of VLM, researchers have started to conceptualize molecules and text as separate modalities, particularly by adopting separate encoders for each modality. As a pioneering work, Text2Mol (Edwards et al., 2021) proposes training separate encoders for molecular graph and textual description with cross-modal contrastive learning. Following this, CLAMP (Seidl et al., 2023) suggests employing contrastive learning for predicting activities based on the textual description of the task. Furthermore, MoleculeSTM (Liu et al., 2023a) develops the largest multi-modal dataset sourced from the PubChem database for cross-modal contrastive learning applications. It is essential to recognize that the distinction between these models is based on their architectural design and the data used for training rather than the training loss itself. On the other hand, MoMu Su et al. (2022) introduces an intermolecular contrastive loss along with random molecular augmentations like node dropping, which may lead to chemically invalid structures Lee et al. (2022a). Unlike previous works, AMOLE concentrates on overcoming the specific hurdles encountered in MoLM: the lack of abundance and expertise in molecule-text pairs, through innovative training loss strategies.
@@
MoLM with Other Multi-Modal Learning. On the other hand, there have been other approaches for integrating molecule text through a novel model architecture. GIMLET (Zhao et al., 2023) suggests encoding graph structure and instructional text directly, without separate graph encoding modules, by utilizing generalized position embeddings. Furthermore, inspired by BLIP-2 (Li et al., 2023) in VLM, MolCA (Liu et al., 2023b) introduces the alignment of two modalities through the Q-Former, allowing the language model to understand 2D molecular graphs. Note that these works are not directly related to our research, as we focus on introducing novel training loss designed for molecule-text pair augmentation and expertise transfer.
3 Preliminaries
3.1 Problem Statement
@@
Notations. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent a molecular graph with atoms $\mathcal{V}$ as nodes and the edges $\mathcal{E}$ given by covalent bonds. Moreover, we have a set of textual descriptions $\mathcal{T}=\{t^{1},\ldots,t^{N}\}$ regarding the molecule $\mathcal{G}$ from $N$ different experts, such as FDA and LiverTox, in Figure 1, each of which details various attributes of the molecule. Note that the number of descriptions for each molecule varies, i.e., $N_{i}$ depends on molecule $\mathcal{G}_{i}$.
@@
Task Description. Given a molecule graph $\mathcal{G}$ with its textual description $t\in\mathcal{T}$, we aim to train encoders $f_{\text{mol}}$ and $f_{\text{text}}$ that produce a molecule representation $z_{\mathcal{G}}\in\mathbb{R}^{D}$ and a textual representation $z_{t}\in\mathbb{R}^{D}$, respectively. We aim to obtain the encoders that produce a generalized representation of molecules and text, that can be utilized for a wide range of downstream tasks.
3.2 Tanimoto Similarity
@@
One traditional way of representing a molecule is through fingerprints, a series of binary bits indicating the presence or absence of specific substructures Rogers and Hahn (2010). The Tanimoto similarity is a widely accepted criterion for calculating the similarity between two molecules Bajusz et al. (2015) based on the fingerprints. Specifically, for the molecules $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$ that are represented with the fingerprints $\mathtt{fp}_{i}$ and $\mathtt{fp}_{j}$, respectively, the Tanimoto similarity is calculated as follows:
$s_{ij}=\frac{|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}{|\mathtt{fp}_{i}|+|\mathtt{fp}_{j}|-|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}.$ (1)
@@
3.3 Molecule-Text Contrastive Learning
@@
Previous works (Liu et al., 2023a; Su et al., 2022) have introduced multi-modal contrastive learning to obtain encoders that establish qualified joint space between a molecule $\mathcal{G}$ and its corresponding text $t$. This approach ensures that paired molecules and texts are closely aligned in the representation space, while unpaired ones remain distant. Specifically, given a molecule $\mathcal{G}$ and its corresponding text $t$, we first obtain the molecule and text representations as follows: $z_{\mathcal{G}}=f_{\text{mol}}(\mathcal{G})$ and $z_{t}=f_{\text{text}}(t)$. Then, the model is trained with the following Noise-Contrastive Estimation (InfoNCE) (Oord et al., 2018) loss:
@@
$\begin{split}\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{2}\Bigl{\{}\log\frac{\exp(\mathtt{sim}(z_{\mathcal{G}},z_{t})/\tau)}{\sum_{t^{\prime}}{\exp(\mathtt{sim}(z_{\mathcal{G}},z_{t^{\prime}})/\tau)}}\\ +\log\frac{\exp(\mathtt{sim}(z_{t},z_{\mathcal{G}})/\tau)}{\sum_{\mathcal{G}^{\prime}}{\exp(\mathtt{sim}(z_{t},z_{\mathcal{G}^{\prime}})/\tau)}}\Bigr{\}},\end{split}$ (2)
@@
where $t^{\prime}$ and $\mathcal{G}^{\prime}$ represent all the texts and molecules within the batch, respectively, $\mathtt{sim}(\cdot,\cdot)$ indicates the cosine similarity between the representations, and $\tau$ denotes the temperature hyperparameter.
4 Methodology
@@
In this section, we introduce our proposed method, called AMOLE, a novel molecule-text contrastive learning approach that addresses the two unique challenges faced in MoLM, i.e., scarcity of molecule-text paired data in quantity and expertise. Figure 2 presents the overall model architecture.
Figure 2: Overall model architecture of AMOLE.
4.1 Augmenting Molecule-Text pairs
@@
While image-text pairs can be easily generated by the general public, molecule-text pairs are typically produced by specialized experts, making it more challenging to augment the data with external sources such as Web. To this end, we propose to augment the molecule-text pair by sharing the textual description among molecules within the data. Building on the well-established biochemical principle that structurally similar molecules often display analogous biological activities (Martin et al., 2002), we propose to share textual descriptions among molecules with structural resemblances. Specifically, we begin by computing all the pairwise structural similarity, i.e., Tanimoto similarity (Eq. 1), between the molecules in the training set. Then, given the similarity information of a molecule $\mathcal{G}_{i}$ to other molecules, we identify the top $k$ molecules that exhibit the highest similarity to $\mathcal{G}_{i}$. We represent these molecules as the set $\mathcal{S}_{i}$. During training, a molecule $\mathcal{G}_{i^{\prime}}$ is randomly selected from the set $\mathcal{S}_{i}$ and substituted for the original molecule $\mathcal{G}_{i}$ according to a predetermined probability $p$; otherwise, it is kept identical to the original molecule $\mathcal{G}_{i}$. By doing so, the textual description $t_{i}$ of molecule $\mathcal{G}_{i}$ is shared among $k$ molecules within the set $\mathcal{S}_{i}$, generating a new molecule-text pair ($\mathcal{G}_{i^{\prime}}$, $t_{i}$). As a result, this approach effectively expands the initial $n$ molecule-text pairs to $n\cdot k$ pairs.
4.2 Structural Similarity Preserving Loss

### v3

# Vision Language Model is NOT All You Need: Augmentation Strategies for Molecule Language Models

@@ Abstract.
Recently, there has been a growing interest among researchers in understanding molecules and their textual descriptions through molecule language models (MoLM). However, despite some early promising developments, the advancement of MoLM still trails significantly behind that of vision language models (VLM). This is because unique challenges exist apart from VLM in the field of MoLM due to 1) a limited amount of molecule-text paired data and 2) missing expertise that occurred due to the specialized areas of focus among the experts. To this end, we propose AMOLE, which 1) augments molecule-text pairs with structural similarity preserving loss, and 2) transfers the expertise between the molecules. Specifically, AMOLE enriches molecule-text pairs by sharing descriptions among structurally similar molecules with a novel structural similarity preserving loss. Moreover, we propose an expertise reconstruction loss to transfer knowledge from molecules that have extensive expertise to those with less expertise. Extensive experiments on various downstream tasks demonstrate the superiority of AMOLE in comprehending molecules and their descriptions, highlighting its potential for application in real-world drug discovery. The source code for AMOLE is available at  https://github.com/Namkyeong/AMOLE.
Molecular Science, Language Model, Drug Discovery
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679607isbn: 979-8-4007-0436-9/24/10ccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Unsupervised learningccs: Computing methodologies Natural language processing
1. Introduction
@@
Machine learning (ML) techniques have recently brought significant advancements to molecular science (Zhang et al., 2023; Wang et al., 2023a; Lee et al., 2023), a field that has traditionally relied on theory, experimentation, and computationally intensive simulations. Particularly, Language Models (LMs), inspired by their recent successes in various fields, have rapidly become popular in molecular science, driven by the wealth of literature available in this domain. Notably, the conventional method of representing molecules as strings, such as SMILES strings (Weininger, 1988), facilitates the integration of two different modalities, i.e., text and molecules, into a single LM. Then, following the masked language modeling (Devlin et al., 2018), Zeng et al. (2022) train the model on masked SMILES and text from the scientific literature. Moreover, inspired by T5 models (Raffel et al., 2020), training the Molecule Language Model (MoLM) with multiple tasks and fine-tuning the downstream task has been proposed (Edwards et al., 2022; Pei et al., 2023; Christofidellis et al., 2023). However, all these models require molecules to be represented in a string format, like SMILES, to be understood by language models.
@@
As an alternative approach, one can treat molecules and language as multiple modalities, following the recent success of the vision language model (VLM). For instance, several recent works (Edwards et al., 2021; Liu et al., 2023b; Seidl et al., 2023) have proposed using separate encoders for each modality and training the paired modalities to be similar in the representation space using contrastive learning, especially inspired by the CLIP model (Radford et al., 2021). These models successfully capture complementary information between different modalities by learning a joint representation space that can be utilized for various downstream tasks such as cross-modal retrieval and molecular property prediction.
@@
Figure 1. (a) Rivastigmine’s textual descriptions from various experts. (b) The majority of molecules in the PubChem database have only one description provided by an expert.
@@
Despite promising early strides, the progress of MoLM lags far behind its VLM counterparts due to the scarcity of molecule-text paired data, both in quantity and expertise. First, in terms of quantity, the VLM community largely follows the viewpoint that scale is everything, as image-text pairs are widely available on the web (Wang et al., 2023b). Pre-training VLM on the massive amounts of crawled image-text pairs, ranging from tens of thousands to billions, consistently leads to significant performance gains in various downstream tasks (Jia et al., 2021; Yu et al., 2022). In contrast, MoLM faces a bottleneck due to the limited amount of molecule-text paired data available, since costly domain knowledge and significant time investment in wet lab experiments are required for generating molecule-text pairs. Therefore, although it is typical to expand image-text pairs by crawling the web, augmenting molecule-text pairs is not straightforward.
@@
Second, in terms of expertise, each molecule has various descriptions provided by different experts, each focusing on its unique areas of expertise. For example, in the case of a molecule named Rivastigmine (Figure 1 (a)), FDA explains its function as a cholinesterase inhibitor, whereas LiverTox (Hoofnagle, 2013) focuses on its effectiveness on Alzheimer’s disease. However, due to the specialized areas of the experts, they typically restrict their knowledge to a selective group of molecules, resulting in numerous molecules having missing expertise across various experts. Specifically, our examination of the largest molecule database, PubChem (Kim et al., 2021), as illustrated in Figure 1 (b), reveals that only 27K out of 299K molecules have descriptions from multiple experts. Consequently, the remaining 272K molecules are each documented with a single expert description, which may lack comprehensive expertise about the molecule.
@@
In this paper, we propose AMOLE that addresses the unique challenges faced in MoLM by Augmenting MOLecule-text pair and transferring Expertise between the molecules. To overcome the lack of molecule-text paired data, we utilize the idea that molecules with similar structures have similar properties, which is grounded in the well-established biochemical principle (Martin et al., 2002). Specifically, we propose to augment molecule-text pairs by sharing descriptions among structurally similar molecules. However, naively treating augmented molecule-text pairs as positive pairs may not be beneficial, as the description is not specifically tailored for structurally similar molecules. To address this issue, instead of directly guiding the model to align augmented pairs closely to the representation, we introduce a novel loss function which instructs the model to align the augmented pairs more accurately based on the structural similarity of the molecules sharing the description.
@@
In this paper, we make the following contributions:
@@
To the best of our knowledge, this is the first paper addressing the data scarcity in MoLM in terms of quantity and expertise.
2. Related Works
@@
MoLM with a Single Language Model. Thanks to the wealth of literature and the traditional string-based representation of molecules, such as SMILES, LMs have been applied to the domain of molecular science. Drawing inspiration from the masked language model approach used in BERT training (Devlin et al., 2018), KV-PLM (Zeng et al., 2022) proposes to train LMs by reconstructing masked SMILES and textual data. Moreover, MolT5 (Edwards et al., 2022) proposes to pre-train the LMs with the “replace corrupted spans” objective (Raffel et al., 2020) on both SMILES string and textual data, followed by fine-tuning for tasks such as molecule captioning and generation. Pei et al. (2023) and Christofidellis et al. (2023) extend MolT5 with various pre-training tasks, such as protein FASTA reconstruction and chemical reaction prediction. However, these models rely on string-based representations of molecules, e.g., SMILES, which are broadly recognized for their lack of topology awareness (Rong et al., 2020). Furthermore, merging two modalities into a single model prevents the adoption of existing superior pre-trained models tailored for each modality (Liu et al., 2023b).
@@
MoLM with Multi-Modal Contrastive Learning. Inspired by the recent success of VLM, researchers have started to conceptualize molecules and text as separate modalities, particularly by adopting separate encoders for each modality. As a pioneering work, Text2Mol (Edwards et al., 2021) proposes training separate encoders for molecular graph and textual description with cross-modal contrastive learning. Following this, CLAMP (Seidl et al., 2023) suggests employing contrastive learning for predicting activities based on the textual description of the task. Furthermore, MoleculeSTM (Liu et al., 2023b) develops the largest multi-modal dataset sourced from the PubChem database for cross-modal contrastive learning applications. It is essential to recognize that the distinction between these models is based on their architectural design and the data used for training rather than the training loss itself. On the other hand, MoMu (Su et al., 2022) introduces an intermolecular contrastive loss along with random molecular augmentations like node dropping, which may lead to chemically invalid structures (Lee et al., 2022a). Unlike previous works, AMOLE concentrates on overcoming the specific hurdles uniquely encountered in MoLM: the lack of abundance and expertise in molecule-text pairs, through innovative training loss strategies.
@@
MoLM with Other Multi-Modal Learning. While the majority of MoLM relies on CLIP style architecture (Radford et al., 2021), there have been other approaches for integrating molecule text through a novel model architecture. As an example, GIMLET (Zhao et al., 2023) suggests encoding graph structure and instructional text directly, without separate graph encoding modules, by utilizing generalized position embeddings. Additionally, drawing inspiration from BLIP-2 (Li et al., 2023) in VLM, MolCA (Liu et al., 2023a) implements a three-stage training pipeline. The initial stage focuses on training the Q-Former to extract molecule representations relevant to the text, and the second stage targets aligning molecular graphs with texts through language modeling. The final stage involves fine-tuning the model for downstream generation tasks. Note that these works are not directly related to our research, as we focus on introducing novel training loss designed for molecule-text pair augmentation and expertise transfer focusing on CLIP-style architecture.
3. Preliminaries
3.1. Problem Statement
@@
Notations. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent a molecular graph with atoms $\mathcal{V}$ as nodes and the edges $\mathcal{E}$ given by covalent bonds. Moreover, we have a set of textual descriptions $\mathcal{T}=\{t^{1},\ldots,t^{N}\}$ regarding the molecule $\mathcal{G}$ from $N$ different experts, such as FDA and LiverTox, in Figure 1 (a), each of which details various attributes of the molecule. Note that the number of descriptions for each molecule varies, i.e., $N_{i}$ depends on molecule $\mathcal{G}_{i}$.
@@
Task Description. Given a molecular graph $\mathcal{G}$ with its textual description $t\in\mathcal{T}$, we aim to train encoders $f_{\text{mol}}$ and $f_{\text{text}}$ that produce a molecule representation $z_{\mathcal{G}}\in\mathbb{R}^{D}$ and a textual representation $z_{t}\in\mathbb{R}^{D}$, respectively. More specifically, we aim to obtain the encoders that produce a generalized representation of molecules and text, that can be utilized for a wide range of downstream tasks, such as zero-shot cross-modal retrieval and zero-shot virtual screening.
3.2. Tanimoto Similarity
@@
One traditional way of representing a molecule is through fingerprints, a series of binary bits indicating the presence or absence of specific substructures (Rogers and Hahn, 2010). The Tanimoto similarity is a widely accepted criterion for calculating the similarity between two molecules (Bajusz et al., 2015) based on the fingerprints. Specifically, for the molecules $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$ that are represented with the fingerprints $\mathtt{fp}_{i}$ and $\mathtt{fp}_{j}$, respectively, the Tanimoto similarity is calculated as follows:
(1) $s_{ij}=\frac{|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}{|\mathtt{fp}_{i}|+|\mathtt{fp}_{j}|-|\mathtt{fp}_{i}\cap\mathtt{fp}_{j}|}.$
@@
3.3. Molecule-Text Contrastive Learning
@@
Previous works (Liu et al., 2023b; Su et al., 2022) have introduced multi-modal contrastive learning to obtain encoders that establish qualified joint space between a molecule $\mathcal{G}$ and its corresponding text $t$. This approach ensures that paired molecules and texts are closely aligned in the representation space, while unpaired ones remain distant. Specifically, given a molecule $\mathcal{G}$ and its corresponding text $t$, we first obtain the molecule and text representations as follows: $z_{\mathcal{G}}=f_{\text{mol}}(\mathcal{G})$ and $z_{t}=f_{\text{text}}(t)$. Then, the model is trained with the following Noise-Contrastive Estimation (InfoNCE) (Oord et al., 2018) loss:
@@
(2) $\begin{split}\mathcal{L}_{\text{InfoNCE}}=-\frac{1}{2}\Bigl{\{}\log\frac{\exp(\mathtt{sim}(z_{\mathcal{G}},z_{t})/\tau)}{\sum_{t^{\prime}}{\exp(\mathtt{sim}(z_{\mathcal{G}},z_{t^{\prime}})/\tau)}}\\ +\log\frac{\exp(\mathtt{sim}(z_{t},z_{\mathcal{G}})/\tau)}{\sum_{\mathcal{G}^{\prime}}{\exp(\mathtt{sim}(z_{t},z_{\mathcal{G}^{\prime}})/\tau)}}\Bigr{\}},\end{split}$
@@
where $t^{\prime}$ and $\mathcal{G}^{\prime}$ represent all the texts and molecules within the batch, respectively, $\mathtt{sim}(\cdot,\cdot)$ indicates the cosine similarity between the representations, and $\tau$ denotes the temperature hyperparameter, which controls the strength of penalties on negative samples (Wang and Liu, 2021).
4. Methodology
@@
In this section, we introduce our proposed method, called AMOLE, a novel molecule-text contrastive learning approach that addresses the two unique challenges faced in MoLM, i.e., scarcity of molecule-text paired data in terms of quantity and expertise. We first explain how to increase the number of molecule-text pairs (Section 4.1). Then, we introduce our novel loss function that aligns augmented pairs more precisely in the representation space (Section 4.2). Finally, we show how to transfer the expertise between molecules (Section 4.3). Figure 2 presents the overall model architecture of AMOLE.
4.1. Augmenting Molecule-Text pairs
@@
In the VLM community, it is well-recognized that a substantial quantity of image-text pairs is crucial for training the model (Jia et al., 2021; Yu et al., 2022), which can be easily accomplished by crawling the image-text pairs from the Web. However, molecule-text pairs are typically produced by specialized experts, making it more challenging to augment the data with external sources such as Web. To this end, we propose to augment the molecule-text pair by sharing the textual description among molecules within the existing data. The key idea here is to share textual descriptions among molecules with structural resemblances, based on the well-established biochemical principle that structurally similar molecules often display analogous biological activities (Martin et al., 2002). In other words, we posit that molecules with similar structures can share textual descriptions that detail their biochemical properties.
Specifically, we begin by computing all the pairwise structural similarity, i.e., Tanimoto similarity (Eq. 1), between the molecules in the training set. Then, given the similarity information of a molecule $\mathcal{G}_{i}$ to other molecules, we identify the top $k$ molecules that exhibit the highest similarity to $\mathcal{G}_{i}$. We represent these molecules as the set $\mathcal{S}_{i}$. During training, a molecule $\mathcal{G}_{i^{\prime}}$ is randomly selected from the set $\mathcal{S}_{i}$ and substituted for the original molecule $\mathcal{G}_{i}$ according to a predetermined probability $p$; otherwise, it is kept identical to the original molecule $\mathcal{G}_{i}$. By doing so, the textual description $t_{i}$ of molecule $\mathcal{G}_{i}$ is shared among structurally similar $k$ molecules within the set $\mathcal{S}_{i}$, generating a new molecule-text pair ($\mathcal{G}_{i^{\prime}}$, $t_{i}$). As a result, this approach effectively expands the initial $n$ molecule-text pairs to $n\cdot k$ pairs without costly wet-lab experiments or the expense of labeling with domain expertise. Note that $\mathcal{S}_{i}$ for each molecule $\mathcal{G}_{i}$ can be readily defined before the model training, thus do not require any additional computational cost during the training process.
Figure 2. Overall model architecture of AMOLE.
4.2. Structural Similarity Preserving Loss